You are 36 miles from a friend. You both start riding your bicycles toward eachother at the same time. You travel 15 mph and she 3 mph slower. How far will you travel befor eyou see your friend?
I need a let statement for x and y

Answer by mananth(16949)

you are 36 miles apart
your speed =15 mph
her speed = 12 mph
..
travelling towards each other so add up the speeds
15+12=27 mph
...
t= distance/speed
t=36/27
t= 4/3 hours
...
time = 4/3
your speed = 15 mph
d= rt
d= 15*4/3
d= 5*4
d=20 miles .
you will meet her after you ride 20 miles.
...
m.ananth@hotmail.ca